### v1

@@ Appendix A Proof of Theorem 1
[…]
$\lim_{N\to\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq\beta_{\min}\cdot\sigma.$ (18)
In particular, when $\beta_{\min}\to 0$,
$\lim_{N\to\infty}\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\leq 0,$ (19)
which implies
$\mathbb{E}\bigg[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\bigg]\xrightarrow{N\to\infty}0.$ (20)
Appendix B Experimental Details
B.1 Environments
The MuJoCo environments used in our experiments (Section 5) follow the same configuration as QARL (Reddi et al., 2024), with the magnitude of the adversarial force tailored to each environment. To avoid catastrophic policy collapse, the maximum adversarial force for each environment is explicitly capped. All environment-specific parameters are listed in Table 3. Note that the adversary action spaces reported in Table 3 do not apply to CAT (Sheng et al., 2022) and MixedNE-LD (Kamalaruban et al., 2020), as in these methods the protagonist and adversary share the same action space.
B.2 Hyperparameter Setting
   Hyperparameter    Value
   Shared (ALL)
    # iterations NN    200
    # episodes per agent per iteration    5
    # evaluation rollouts per iteration    10
    # hidden layers    3
    # hidden units per layer    256
    discount factor    0.99
    horizon    500
   UACER
    # ensemble Critic networks $K$    5
    $\beta_{0}$    0.85
    $\beta_{min}$    0.15
    decay speed $\lambda$    3
   Shared (SAC and SAC MixedNE-LD)
    nonlinearity    ReLU
    critic optimiser    Adam
    critic learning rate    $3\times 10^{-4}$
    actor learning rate    $1\times 10^{-4}$
    initial replay memory size    $3\times 10^{3}$
    max replay memory size    $1\times 10^{6}$
    warmup transitions    $5\times 10^{3}$
    batch size    256
    target smoothing coefficient ($\tau$)    $5\times 10^{-3}$
    target update interval    1
    policy log std bounds    [20,2][-20,2]
    initial temperature    $5\times 10^{-3}$
    temperature learning rate    $3\times 10^{-4}$
    target entropy    $\textrm{dim}(\mathcal{A})$
   SAC
    actor optimiser    Adam
   SAC MixedNE-LD
    adversary influence $\sigma$    0.1
    actor optimiser    SGLD
    thermal noise ($\sigma_{t}$)    $10^{-3}\times(1-5\times 10^{-5})^{t}$
    RMSProp parameter $\alpha$    0.999
    RMSProp parameter $\epsilon$    $10^{-8}$
   CAT
    curriculum start iteration    $0.2\times\text{\# iterations}$
    curriculum end iteration    $0.8\times\text{\# iterations}$
   CAT Adversary
    gradient descent learning rate    3
    gradient descent step limit    25
    gradient descent convergence threshold $\epsilon$    $10^{-3}$
    disturbance $L^{p}$-norm    2
   QARL
    initial gamma distribution concentration $k_{\text{initial}}$    50
    target gamma distribution concentration $k_{\text{target}}$    1
    fixed gamma distribution rate    1000
    $D_{\text{KL}}$ constraint ($\epsilon$)    0.5
    # rollouts needed for update    30
Table 4: Algorithm hyperparameters.
To ensure fair and valid comparisons, we strictly follow the experimental protocol established by QARL (Reddi et al., 2024), maintaining consistent shared hyperparameter settings across all methods. This practice ensures that performance differences can be attributed to algorithmic distinctions rather than variations in parameter tuning. The hyperparameters used for UACER and each baseline method are provided in Table 4.
B.3 Hardware and Software
All experiments were run on a workstation equipped with an NVIDIA TITAN Xp GPU (12 GB), 15 GB of system memory, and an Intel® Xeon® E5-2680 v4 CPU. The implementations of the algorithms are based on the publicly available codebase of Reddi et al. (2024), which itself is built on the MushroomRL library (D’Eramo et al., 2021). The same library is also used to implement the agents and to construct the adversarial environment wrappers.

### v2

@@ IV-C Theoretical Guarantees for TDU
$\lim_{N\rightarrow\infty}E$ $\Big[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\Big]$ (19)
$\leq$ $\lim_{N\rightarrow\infty}E\Big[\big|\mu_{Q}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\Big]$
$+\beta(n)\cdot\lim_{N\rightarrow\infty}E\big[\sigma_{Q}(s_{t},a_{t})\big]$
$=$ $\;\beta_{\min}\cdot\lim_{N\rightarrow\infty}E\big[\sigma_{Q}(s_{t},a_{t})\big].$
$\lim_{N\to\infty}\mathbb{E}\Big[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\Big]\leq\beta_{\min}\cdot\sigma.$ (20)
In particular, when $\beta_{\min}\to 0$,
$\lim_{N\to\infty}\mathbb{E}\Big[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\Big]\leq 0,$ (21)
which implies
$\mathbb{E}\Big[\big|Q_{E}(s_{t},a_{t})-Q^{\ast}(s_{t},a_{t})\big|\Big]\xrightarrow{N\to\infty}0.$ (22)
V Experimental Results
  • Q1
    Can UACER successfully achieve more robust (see Figs. 4, 5 and Table IV), stable (see Table III), and efficient (see Fig. 4) adversarial RL compared with existing methods?
  • Q2
    How crucial is each proposed component in UACER for improving RL performance (see Fig. 6)?
  • Q3
    Does the optimistic design of TDU provide measurable benefits in robustness and convergence compared to conventional pessimistic value estimation under adversarial conditions (see Fig. 7)?
  • Q4
    How does critic ensemble size $K$ affect the performance (see Fig. 9)?
  • Q5
    Can UACER be universally applicable to other existing robust adversarial RL algorithms (see Fig. 8)?
V-A Environments
We evaluate UACER on several challenging MuJoCo control tasks [35] from the DeepMind Control Suite [36], including one sparse-reward task, Cartpole-Swingup Sparse, and three complex locomotion tasks, Walker-Run, Hopper-Hop, and Cheetah-Run. As shown in Fig. 3, each task name is followed by two integers that specify the dimensions of the state and action spaces. Following established adversarial training protocols [28, 17, 30], we instantiate two agents: a protagonist, which learns an optimal policy to accomplish the task, and an adversary, which applies external perturbation forces to disrupt the protagonist’s performance, thereby eliciting robust behavior. Both agents observe the standard environmental states as input. The action space of the protagonist coincides with the original action space of the environment, whereas the adversary’s action space is defined as a bounded, multi-dimensional disturbance force. For each environment, the adversary’s perturbation intensity is carefully tuned to be sufficiently large to beget agent robustness and generalization while still posing a meaningful challenge to the protagonist. The MuJoCo environments follow the same configuration as in QARL [30], with the magnitude of the adversarial force customized for each task. To prevent catastrophic policy collapse, the maximum adversarial force is explicitly capped in each environment. All environment-specific adversary parameters are listed in Table II.
Figure 3: Representative challenging MuJoCo control tasks from the DeepMind Control Suite including: (a) Cartpole-Swingup Sparse, $|\mathcal{S}|=4$, $|\mathcal{A}|=1$; (b) Walker-Run, $|\mathcal{S}|=18$, $|\mathcal{A}|=6$; (c) Hopper-Hop, $|\mathcal{S}|=14$, $|\mathcal{A}|=4$; (d) Cheetah-Run, $|\mathcal{S}|=18$, $|\mathcal{A}|=6$.
TABLE II: Environment-specific parameters for the adversary.
Env.
Adversary
Max Force
Performance
Lower Bound
Adversary Action
Description ($|\mathcal{A}_{a}|$)
Cartpole 0.005 10 2D force on the pole (2)
Walker 1.0 10 2D forces on the feet (4)
Hopper 1.0 5 2D forces on the foot & torso (4)
Cheetah 1.0 40 2D forces on the feet & torso (6)
V-B Baselines
We evaluate our approach in comparison to the classical robust adversarial RL method RARL [28] and several recently proposed improvements addressing RARL’s saddle point optimization challenges, including MixedNE-LD [17], which utilizes Langevin dynamics to escape local optima, and Curriculum Adversarial Training (CAT) [31] featuring dynamic adversary strength adjustment via a hand-designed curriculum. Additionally, we compare against the state-of-the-art QARL algorithm [30], which synergistically combines entropy regularization with quantal response equilibrium while employing temperature-based curriculum learning to regulate adversary rationality. For comprehensive assessment, standard SAC [15] serves as a non-adversarial baseline representing the underlying RL framework.
Unless specified otherwise, UACER implementations in the following build upon the QARL framework. For each comparison method, we conduct training over five random seeds $\{0,1,2,3,4\}$ across all environments, with $200$ alternating iterations per seed. All reported results represent the average performance over these five runs. Complete hyperparameter configurations for all algorithms are provided in Appendix.
All experiments were run on a workstation equipped with an NVIDIA TITAN Xp GPU (12 GB), 15 GB of system memory, and an Intel® Xeon® E5-2680 v4 CPU. The implementations of the algorithms are based on the publicly available codebase of QARL [30], which itself is built on the MushroomRL library [9]. The same library is also used to implement the agents and to construct the adversarial environment wrappers.
V-C Evaluation Protocols
To comprehensively assess the effectiveness of the proposed method, we evaluate the learned protagonist from multiple complementary perspectives during and after training. In particular, our evaluation focuses on three key dimensions: the agent’s ability to withstand disturbances, the stability of its learning dynamics, and its final robustness under adversarial conditions. First, to evaluate disturbance resistance during training, we conduct a robustness evaluation every five alternating iterations. For each environment, we sweep through a predefined range of environmental parameter variations (e.g., mass and friction coefficient) in the absence of the adversary, and then report the average return across all parameter configurations as the robustness metric. This procedure introduces graded disturbances via variations in dynamic parameters, offering a thorough test of robustness while simultaneously assessing generalization capability under domain shift conditions. Concurrently, we quantify training stability by computing the percentage decrease in the protagonist’s performance between consecutive (i.e., $i$ and $i+1$) robustness evaluations and averaging these decreases over the entire training process. Finally, following the practice in [30], we evaluate the final adversarial performance of the protagonist against the worst adversary in a minimax sense by fixing the protagonist obtained at the end of training and subsequently training a dedicated adversary against it, and then using that adversary for testing.
Figure 5: Heatmap of final robustness evaluation for UACER and baseline methods on MuJoCo swingup and locomotion tasks, evaluated at the end of training. Each heatmap illustrates the performance achieved for different environmental properties, which are described on the x-y axes.
V-D Results
To examine the effectiveness of our proposed method (in response to Q1), we conduct experiments on the aforementioned MuJoCo control tasks with UACER and all baseline methods. The experimental results and corresponding analyses are presented below.
Robustness and Convergence
As demonstrated by the robustness evaluation results in Fig. 4, UACER achieves substantially faster convergence compared to all baseline methods across most environments, effectively mitigating the long-standing challenge of slow convergence in robust adversarial RL. This framework consistently attains superior robustness scores in all tested tasks, reflecting both exceptional perturbed resilience and generalization ability. Notably, in more complex and sparse-reward environments such as Hopper-Hop and Cartpole-Swingup Sparse, the baselines exhibit noticeably poorer average performance, as the task complexity and sparse reward setting enable the adversary to easily hinder the protagonist’s policy learning. In contrast, UACER achieves more statistically pronounced performance improvements over competing approaches. The key to this success lies in its two core design aspects: the parallel ensemble of value networks reduces Q-value estimation variance and stabilizes policy updates under adversarial perturbations; at the same time, the time-decaying uncertainty-weighted aggregation mechanism adaptively balances exploration and exploitation. By encouraging thorough exploration early in training, it enables the agent to quickly adapt to perturbed environments while maintaining stability throughout the learning process.
Moreover, we visualize the robustness of the trained protagonist to different test conditions upon training completion, as illustrated in Fig. 5. Each heatmap depicts the performance obtained under varying properties of the environment. These results further validate UACER’s capability to handle diverse environmental dynamics and also indirectly indicate that the observed comprehensive robustness gains stem from holistic adaptation to various environmental properties, rather than overfitting to any specific environmental dynamics.
TABLE III: Training stability analysis of UACER and baseline methods on MuJoCo swingup and locomotion tasks (based on the results in Fig. 4). Performance degradation (percentage decrease, $\%$) is reported as the mean $\pm$ standard error over five independent
trials. Values shown in bold and underlined indicate the best and second-best results, respectively.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
SAC $60.504\pm 53.811$ $353\pm 1.862 33.$$3.367\pm 8.704$ $4.520\pm 4.724$
CAT $56.902\pm 51.160$ $4.743\pm 2.955$ $32.276\pm 12.829$ $5.353\pm 3.882$
QARL $35.714\pm 28.414$ $4.527\pm 0.847$ $35.581\pm 6.025$ $8.692\pm 4.927$
UACER $17.352\pm 9.593$ $1.659\pm 0.604$ $27.938\pm 8.474$ $7.665\pm 3.515$
TABLE IV: Final adversarial performance of UACER and baseline methods on MuJoCo swingup and locomotion tasks, evaluated at the
end of training against an adversary trained against the frozen trained protagonist. Performance metric (return) is
reported as the mean $\pm$ standard error over five independent trials. Values shown in bold and underlined indicate
the best and second-best results, respectively. Notably, SAC is omitted as it lacks adversarial components.
Method Cartpole-Swingup Sparse Walker-Run Hopper-Hop Cheetah-Run
UACER $334.770\pm 6.886$ $341.995\pm 8.480$ $96.564\pm 17.479$ $312.594\pm 30.422$
Training Stability
We evaluate the training stability of UACER from two complementary perspectives: the performance variability across different random seeds and the magnitude of return degradation fluctuations during training. As illustrated in Fig. 4, UACER exhibits the smallest variance in performance curves across different random seeds, as evident from the narrowest shaded regions, indicating minimal sensitivity to random initialization. This enhanced stability to initialization can be attributed to the critic ensemble framework, which aggregates Q-value predictions from multiple parallel critics to produce more accurate and stable value estimates, thereby mitigating divergent updates caused by different initialization conditions. Furthermore, the training stability metrics derived from the results in Fig. 4 are summarized in Table III, where performance degradation (percentage decrease, $\%$) is reported as the mean $\pm$ standard error over five independent trials. Values shown in bold and underlined indicate the best and second-best results, respectively. The results demonstrate that UACER achieves the highest training stability in the first three environments, while maintaining performance comparable to the baselines in the last environment (i.e., Cheetah-Run). These findings further validate the effectiveness of UACER in alleviating training fluctuations induced by adversarial disturbances in robust adversarial RL settings.
Worst-Case Adversarial Performance
In addition to the aforementioned comprehensive robustness evaluation under varying environmental parameters, we further follow the minimax principle to examine UACER’s ability to cope with stronger and previously unseen adversaries. Specifically, after training is completed, we fix the learned protagonist and continue to train an adversary to minimize the protagonist’s performance, and then evaluate the final adversarial performance between the protagonist and this subsequently optimized, strongest adversary. As quantitatively summarized in Table IV, UACER consistently outperforms the baseline methods across all environments when facing such further optimized adversaries, while also exhibiting relatively low variance across different initialization settings. These empirical results demonstrate its strong capability to withstand extreme adversarial attacks as well as its superior generalization ability in the presence of untrained perturbations.
Figure 7: Robustness evaluation curves of UACER (optimistic TDU aggregation) and its three pessimistic Q-value estimation variants (UACER with pessimism-dec/pessimism-dec/min aggregation) on MuJoCo swingup and locomotion tasks.
V-E Ablation Study
In response to Q2, we conduct a systematic ablation study to investigate the individual contributions of UACER’s core components, including the critic ensemble, the TDU aggregation mechanism, and the diversity enhancement strategies. Specifically, we perform controlled single-variable ablation experiments in robust adversarial RL settings by comparing UACER with the following three algorithmic variants: (i) w/o ensemble, which replaces the $K$-critic ensemble with the standard two-critic architecture used in the base RL algorithm SAC; (ii) w/o TDU, which substitutes the TDU aggregation function with minimum-value selection strategy adopted in SAC; (iii) w/o diversity, which removes the diversity enhancement strategies from UACER.
The results presented in Fig. 6 indicate that removing any individual component from UACER leads to varying degrees of performance degradation. Specifically, when the diversity enhancement strategy is omitted (“w/o diversity”), the critic networks tend to converge toward similar update patterns, which weakens the ensemble framework’s ability to capture uncertainty under adversarial perturbations. Consequently, the algorithm exhibits noticeably poorer average performance in certain environments, such as Cartpole-Swingup Sparse and Hopper-Hop. More critically, removing either the critic ensemble (“w/o ensemble”) or the TDU aggregation mechanism (“w/o TDU”) results in substantial declines in both robustness and convergence speed across most tasks, particularly in challenging environments characterized by high complexity and sparse rewards. Furthermore, compared with the other two components, eliminating the TDU mechanism has the most pronounced negative impact on performance across all environments, highlighting the crucial role of the proposed uncertainty-adaptive TDU aggregation mechanism.
In contrast, the complete UACER framework, which integrates all components, consistently outperforms its ablated variants. These findings suggest that in environments subject to persistent adversarial disturbances, no single component alone is sufficient to achieve comprehensive robustness across all tasks. Instead, the synergistic integration of all components is essential: the critic ensemble provides a solid foundation for stable value estimation; the TDU mechanism effectively converts base model prediction uncertainty into adaptive and stable policy update signals; and, building upon this foundation, the diversity-enhancing network initialization and architectural design further improve training stability and resistance to adversarial perturbations.
V-F Analysis
Figure 8: Robustness evaluation curves of UACER combined with other existing robust adversarial RL algorithms including RARL and SAC MixedNE-LD on MuJoCo swingup and locomotion tasks.
Figure 9: Robustness evaluation of UACER with varying critic ensemble sizes $K$ on the Walker-Run and Hopper-Hop tasks.
Superiority of TDU’s Optimistic Aggregation in Adversarial RL
@@ V-F Analysis
To evaluate the superiority of TDU’s optimistic aggregation in adversarial RL (in response to Q3), we compare it with three pessimistic variants derived from Eq. (8): (i) pessimism-dec: the mean minus the uncertainty term, with $\beta(n)$ decaying from 1 to 0; (ii) pessimism-inc: the mean minus the uncertainty term, with $\beta(n)$ increasing from 0 to 1; (iii) pessimism-min: the minimum among the $K$ critic outputs. As demonstrated in Fig. 7, TDU consistently delivers superior performance across all tasks. This advantage arises because, in non-stationary adversarial environments, the protagonist already struggles to learn appropriate responses to perturbations, and pessimistic estimation further constrains its Q-value prediction capacity under such disturbances. By contrast, TDU’s optimistic aggregation achieves a more effective balance between bias control and robustness enhancement, thereby alleviating the tendency toward over-conservative policy learning while simultaneously strengthening training stability and final performance.
Effects of Critic Ensemble Size
The ensemble size $K$, i.e., the number of critic networks, is a key hyperparameter newly introduced in the proposed UACER framework. To analyze its impact (in response to Q4), we conduct controlled experiments in which $K$ is varied from $2$ to $10$ while all other hyperparameters remain fixed. Fig. 9 presents box plots illustrating two distinct trends of agent robustness performance with respect to the value of $K$ on the tasks Walker-Run and Hopper-Hop. It can be observed that the agent’s performance does not increase monotonically with larger ensemble sizes, contrary to intuitive expectations. Overall, when $K$ is relatively small (e.g., $K<5$), the limited ensemble diversity is insufficient to fully exploit the stability gains of critic ensemble for Q-value estimation. In this regime, increasing $K$ gradually improves the agent’s robustness performance. However, once $K$ exceeds a certain threshold (e.g., $K>5$), the performance gains begin to saturate and may even decline to varying degrees. A possible explanation for this phenomenon is that, although a larger ensemble size can theoretically enhance diversity and thereby stabilize Q-value estimation, an excessive number of critics may introduce detrimental redundancy and complicate policy optimization. Consequently, a larger $K$ does not necessarily lead to better performance, and its appropriate value should be determined according to the complexity of the specific task. In this work, balancing the agent’s performance across the examined range of $K$ values and the associated computational cost, we select $K=5$ for all experimental validation.
Scalability Analysis
All previous experiments implemented UACER based on the QARL framework as the foundational robust adversarial RL approach. In response to Q5, this section examines UACER’s applicability as a general framework when integrated with other existing robust adversarial RL algorithms. Specifically, we migrate UACER to the classic RARL algorithm and its improved extension SAC MixedNE-LD for practical implementation and evaluation. The learning curves in terms of robustness performance are shown in Fig. 8. The experimental results confirm that both RARL and SAC MixedNE-LD achieve substantial performance gains after incorporating UACER, showing improved convergence speed, training stability, and robustness to adversarial perturbations. In essence, UACER is designed as an algorithm-agnostic framework that is compatible with any value-based robust adversarial RL methods. Regrettably, in the Cartpole-Swingup Sparse task, SAC MixedNE-LD exhibited critically poor baseline performance (near-zero returns throughout training), with no discernible improvement after UACER integration. This outcome indicates that UACER’s performance improvements remain partly dependent on the base adversarial RL algorithm’s intrinsic adaptability to the specific task.
Computational Overhead and Limitations
VI Conclusion
In this article, we present UACER, a novel robust adversarial RL approach that advances the field through two core contributions: (i) a parallel ensemble of critic networks that simultaneously reduces Q-value estimation variance and enhances robustness under adversarial perturbations, and (ii) a time-varying decay uncertainty Q-value aggregation mechanism that adaptively balances exploration and exploitation through variance-based epistemic uncertainty quantification across critics. Supported by theoretical analysis and extensive experimental validation, UACER achieves significant improvements in training stability, convergence speed, and adversarial robustness, effectively addressing critical challenges that hinder the reliable deployment of deep RL systems in real-world applications. By integrating ensemble-based value estimation with uncertainty-adaptive policy optimization, the proposed framework provides a systematic methodology for developing RL algorithms with provable stability guarantees and consistent performance in complex and uncertain environments.
Future research directions include extending the UACER architecture to partially observable settings, where information constraints introduce additional sources of uncertainty, and exploring its applications in real-world safety-critical systems, such as robotic manipulation and autonomous driving scenarios, where both robustness and adaptability are essential for safe and reliable operation.